whats the value of the standard deviation of this normal curve.
0 100 200 300 400 500
Answer #1

From the normal curve the mean is approximately

mu=250

According to the empirical rule 99.7% observations lies within 3 standard deviation of mean.

Therefore, mu+3sigma=450

250+3sigma=450

3sigma=450-250

3sigma=200

Sigma=200/3

Sigma=66.67

Hence the approximate standard deviation is

Sigma=66.67